A mass weighing 39.5 kg. stretches a spring 1/4m. At t=0 , the mass is released from a point 3/4m below the equilibrium position with an upward velocity of (5/4)m/second. Determine the function x(t) that describes the subsequent free motion.
1
Expert's answer
2020-03-13T10:24:21-0400

From Newton's second law "m\ddot{x}=-kx" where "k" is a spring stiffness, typing "\omega^2=\frac{k}{m}" hence "\omega=\sqrt\frac{k}{m}" . From equilibrium "mg=kl" ,where "l=\frac{1}{4}m" then "k=\frac{mg}{l}" then "\omega=\sqrt\frac{g}{l}\approx6.26s^-1"

"x(t)=a\cos(\omega t-\alpha)" ,"v(t)=x_t'=-a\omega\sin(\omega t-\alpha)"

"x_0(0)=a\cos\alpha" ,"v_0(0)=a\omega\sin\alpha" , since "x_0=0.75m" , "v_0=1.25\frac{m}{s}" hence "a=\sqrt{x_0^2+\frac{v_0^2}{\omega^2}}=0.775" m

"\frac{v_0}{x_0}=\frac{a\omega \sin\alpha}{a\cos\alpha}=\omega\tan\alpha" ,hence "\tan\alpha=\frac{v_0}{x_0 \omega}=0.27", hence "\alpha=\arctan0.27\approx15^0\approx0.083\pi" then

"x(t)=0.775\cos(6.26t+0.083)m"